OpenStudy (anonymous):

if f(x) =x-1/x=1 then f^-1(x)=

OpenStudy (anonymous):

\[f(x)=\frac{x+1}{x-1}\]?

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

and you want the inverse right?

OpenStudy (anonymous):

yeh @satellite73

OpenStudy (anonymous):

usual method is to write \[x=\frac{y+1}{y-1}\] and solve for \(y\)

OpenStudy (anonymous):

x-1/x+1 yeh

OpenStudy (anonymous):

takes a few steps to do it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh let me get it right \[f(x)=\frac{x-1}{x+1}\] is the original function yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

not the one i wrote above' ok we start with \[x=\frac{y-1}{y+1}\] and solve for \(y\)

OpenStudy (anonymous):

first step is to multiply by \(y+1\) to get rid of the fraction and write \[x(y+1)=y-1\]

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

next step is to distribute \[xy+x=y-1\]

OpenStudy (anonymous):

then since you want \(y\) by itself, you have to put it on one side of the equal sign say the left, by writing \[xy+x-y=-1\] then put the \(x\) on the right via \[xy-y=-1-x\]

OpenStudy (anonymous):

finally factor the \(y\) out of the left hand side, so you know what to divide by \[y(x-1)=-1-x\]

OpenStudy (anonymous):

and then divide to get \[y=\frac{-1-x}{x-1}\]

OpenStudy (anonymous):

now there are way to many minus signs, so it might be nice to clean it up and have \[f^{-1}(x)=\frac{x+1}{1-x}\]

OpenStudy (anonymous):

i don't think i skipped any steps if any are not clear let me know

OpenStudy (anonymous):

no its clear thanks:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

one thing what about the minus sign?

OpenStudy (anonymous):

when we solved, the answer was ugly, it was \[\frac{-1-x}{x-1}\]

OpenStudy (anonymous):

there is really nothing wrong with that, but it has a lot of minus signs in it if you multiply the top and bottom by \(-1\) i.e. change all the signs, you get the nicer looking \[\frac{x+1}{1-x}\]

OpenStudy (anonymous):

i.e. \[\frac{-1-x}{x-1}\times \frac{-1}{-1}=\frac{x+1}{1-x}\]

OpenStudy (anonymous):

ok thanks now got that

OpenStudy (anonymous):

yw again